OpenStudy (anonymous):

"A laser is used in studies of initiating controlled nuclear fusion . It can deliver a power of 1.60 × 1013 W over a time interval of 2.50 ns." I need to find the energy output. so I just thought about how power (Watts) is Joules over a specific time interval.... so if the Power is 1.60 x 1013 Watts over a 2.50 ns interval, and P=J/Δt, then if I rearrange the equation to 1.60 x 1013 W * 2.50 x 10-9 s = Joules, i would also get the joules/second (aka Watts with a time interval of 1 s) because Joules over 1 second is the same number... but i'm not sure if i'm correct.

OpenStudy (michele_laino):

\(energy= power \cdot time \)

OpenStudy (michele_laino):

so we get: \[W = 1.6 \cdot {10^{13}} \cdot 2.5 \cdot {10^{ - 9}} = 4.0 \cdot {10^4}Joules\]

OpenStudy (michele_laino):

\(W \) stands for energy